### v1

# Oriented discrepancy of Hamilton cycles and paths in digraphs

Abstract
Erdős (1963) initiated extensive graph discrepancy research on 2-edge-colored graphs. Gishboliner, Krivelevich, and Michaeli (2023) launched similar research on oriented graphs. They conjectured the following generalization of Dirac’s theorem: If the minimum degree $\delta$ of an $n$-vertex oriented graph $G$ is greater or equal to $n/2$, then $G$ has a Hamilton oriented cycle with at least $\delta$ forward arcs. This conjecture was proved by Freschi and Lo (2024) who posed an open problem to extend their result to an Ore-type condition. We propose two conjectures for such extensions and prove some results which provide support to the conjectures. For forward arc maximization on Hamilton oriented cycles and paths in semicomplete multipartite digraphs and locally semicomplete digraphs, we obtain characterizations which lead to polynomial-time algorithms.
Keywords: oriented discrepancy; Hamilton oriented cycles; forward arcs; semicomplete multipartite digraphs
1 Introduction
Combinatorial discrepancy is studied on hypergraphs. For a hypergraph $\cal H$ and coloring $c:V({\cal H})\rightarrow\{-1,1\}$, the discrepancy of an edge $e\in E({\cal H})$ is $D(e)=|\sum_{v\in e}c(v)|.$ Combinatorial discrepancy studies $\min_{c}\max_{e\in E({\cal H})}D(e)$, the discrepancy of $\cal H$; for an excellent exposition of the topic, see e.g. [23]. Erdős [12] launched an extensive study of graph discrepancy, which is a special case of combinatorial discrepancy: $V({\cal H})$ is the edge set of some graph $G$ and $E({\cal H})$ consists of the edge sets of specific subgraphs of $G.$ For recent papers on the topic, see e.g. [2, 3, 14, 15, 16]. In particular, Balogh et al. [2] proved the following result.
Theorem 1.1.
Let $40<t<1/$ 4 and let $n$ be a sufficiently large positive integer. Then every 2-edge-coloring of an $n$-vertex graph $G$ with minimum degree $\delta(G)\geq(\frac{3}{4}+t)n,$ has a Hamilton cycle with at least $(\frac{1}{2}+\frac{t}{64})n$ edges of the same colour (and so the discrepancy at least $tn/32$).
Motivated by Theorem 1.1, Gishboliner, Krivelevich, and Michaeli [17] introduced the oriented version of graph discrepancy and conjectured the result of Theorem 1.2, mentioned below and proved by Freschi and Lo [13].
Let us now introduce some terminology and notation; terminology and notation not introduced in this paper can be found in [6, 7, 9]. The underlying graph of a digraph $D$ is an undirected multigraph $U(D)$ obtained from $D$ by removing the orientations of all arcs in $D$. The degree of a vertex $x$ in a digraph $D$, denoted by $d_{D}(x)$, is the degree $d_{U(D)}(x)$ of $x$ in $U(D)$. In what follows, we will often omit directed or undirected graph subscripts if such graphs are clear from the context. The minimum degree of a vertex in a directed or undirected graph $H$ is denoted by $\delta(H)$. A digraph $D$ is an oriented graph if there is no more than one arc between any pair of vertices in $D$. An oriented cycle (oriented path, respectively) in $D$ is a subdigraph $Q$ of $D$ such that $U(Q)$ is a cycle (path, respectively) in $U(D)$. An oriented cycle (oriented path, respectively) in $D$ is Hamilton if it is a spanning subdigraph of $D$. Note that a digraph $D$ has a Hamilton oriented cycle (path, respectively) if and only if $U(D)$ has a Hamilton cycle (path, respectively).
Let $C=v_{1}v_{2}\dots v_{p}v_{1}$ be an oriented cycle. The arc between $v_{i}$ and $v_{i+1}\pmod{p}$ on $C$ is forward if $v_{i}v_{i+1}\in A(C)$ and backward if $v_{i+1}v_{i}\in A(C)$. The number of forward (backward, respectively) arcs of $C$ is denoted by $\sigma^{+}(C)$ ($\sigma^{-}(C)$, respectively). If all arcs of $C$ are forward, then $C$ is a directed cycle (or, just a cycle). While $\sigma^{+}(C)$ and $\sigma^{-}(C)$ depend on the order of its vertices, $\sigma_{\min}(C)=\min\{\sigma^{+}(C),\sigma^{-}(C)\}$ and $\sigma_{\max}(C)=\max\{\sigma^{+}(C),\sigma^{-}(C)\}$ do not depend on the order. Similar definitions and notation can be introduced for oriented paths. A digraph $D$ is hamiltonian if it has a Hamilton cycle.
Theorem 1.2 ([13]).
Let $D$ be an oriented graph on $n\geq 3$ vertices. If $\delta(D)\geq\frac{n}{2}$, then there exists a Hamilton oriented cycle $C$ in $D$ such that $\sigma_{\max}(C)\geq\delta(D)$.
Note that Theorem 1.2 is a strengthening of Dirac’s theorem [11]: a graph $G$ on $n\geq 3$ vertices has a Hamilton cycle if $\delta(G)\geq n/2.$ Ore [24] extended Dirac’s theorem as follows: a graph $G$ on $n\geq 3$ vertices is hamiltonian, i.e. has a Hamilton cycle, if $d(x)+d(y)\geq n$ for every pair $x,y$ of non-adjacent vertices of $G$. In a digraph $D$, a pair of vertices are non-adjacent if there is no arc between them.
Freschi and Lo [13] stated an open problem to extend their result to an Ore-type condition. Now we pose two conjectures on the open problem. While we were unable to prove either conjecture, we show some results providing support to the conjectures.
By considering the minimum degree of $D$, we may extend Theorem 1.2 to the following conjecture, which is a natural analogue of Theorem 1.2 with an Ore-type condition.
Conjecture 1.3.
Let $D$ be an oriented graph on $n\geq 3$ vertices with minimum degree $\delta$. If $d(u)+d(v)\geq n$ for each pair of non-adjacent vertices $u$ and $v$, then there exists a Hamilton oriented cycle $C$ in $D$ such that $\sigma_{\max}(C)\geq\max\{\delta,n-\delta\}$.
When $\delta\geq n/2$, Conjecture 1.3 holds and is sharp, as it directly follows from Theorem 1.2. We now assume that $\delta<n/2$. In this case, we need to find a Hamilton oriented cycle $C$ in $D$ such that $\sigma_{\max}(C)\geq n-\delta$. In fact, this bound is sharp for a family of examples that contain no Hamilton oriented cycle $C$ with $\sigma_{\max}(C)>n-\delta$. We will present the examples and prove some results supporting Conjecture 1.3 in Section 2.
For every $n$-vertex oriented graph $D$, we define
$s^{*}(D)=\min\{d(u)+d(v)-n:u\neq v\in V(D),\{uv,vu\}\cap A(D)=\emptyset\}$
if there is a pair of non-adjacent vertices in $D$, and $s^{*}(D)=n-2$ if $D$ is a tournament. The following conjecture is another natural generalization of Theorem 1.2.
Conjecture 1.4.
Let $D$ be an oriented graph on $n\geq 3$ vertices. If $s^{*}(D)\geq 0$, then there is a Hamilton oriented cycle $C$ in $D$ such that $\sigma_{\max}(C)\geq\lceil\frac{n+s^{*}(D)}{2}\rceil$.
If Conjecture 1.4 is true, it also implies Theorem 1.2. Indeed, when $\delta(D)\geq n/2$, since $s^{*}(D)\geq 2\delta(D)-n$, Conjecture 1.4 implies that $\sigma_{\max}(C)\geq n/2+(2\delta(D)-n)/2=\delta(D)$. Furthermore, the bound is tight, as we construct a family of oriented graphs on $n$ vertices that contain no Hamilton oriented cycle $C$ with $\sigma_{\max}(C)>\lceil\frac{n+s^{*}(D)}{2}\rceil$. We will show the corresponding constructions and prove an approximate version of Conjecture 1.4 in Section 2.
A semicomplete digraph is a digraph obtained from a complete graph by replacing every edge $\{x,y\}$ with either arc $xy$ or arc $yx$ or two arcs $xy$ and $yx.$ It is well-known that every semicomplete digraph has a Hamilton path and every strongly connected semicomplete digraph has a Hamilton cycle, see e.g. [7]. These two results make it straightforward to find a Hamilton oriented path (cycle, respectively) with maximum number of forward arcs in a semicomplete digraph $D.$ In Section 3, we consider semicomplete multipartite digraphs, which are generalizations of semicomplete digraphs. A semicomplete $p$-partite (or, multipartite) digraph is a digraph obtained from a complete $p$-partite graph ($p\geq 2$) by replacing every edge $\{x,y\}$ with either arc $xy$ or arc $yx$ or two arcs $xy$ and $yx.$ Maximal independent vertex sets of a semicomplete multipartite digraph are called its partite sets. For a recent survey on paths and cycles in semicomplete multipartite digraphs, see [28]. In Section 3, we prove two theorems which determine the maximum $\sigma_{\max}(Q)$ for a Hamilton oriented path (cycle, respectively) $Q$ in a semicomplete multipartite digraph $D$ provided $D$ has a Hamilton oriented path (cycle, respectively).
Let $D$ be a semicomplete multipartite digraph with partite sets of sizes $n_{1},\dots,n_{p}$. We say that $D$ satisfies the HC-majority inequality ( HP-majority inequality, respectively) if $2\max\{n_{i}:\ i\in[p]\}\leq\sum_{i=1}^{p}n_{i}$ ($2\max\{n_{i}:\ i\in[p]\}\leq(\sum_{i=1}^{p}n_{i})+1,$ respectively). It is not hard to prove that a semicomplete multipartite digraph $D$ with partite sets of sizes $n_{1},\dots,n_{p}$ has a Hamilton oriented cycle (path, respectively) if and only if $D$ satisfies the HC-majority inequality (the HP-majority inequality, respectively).
A 1-path-cycle factor in a digraph $H$ is a spanning subdigraph of $H$ consisting of a path and a collection of cycles, all vertex-disjoint. Note that a Hamilton path is a 1-path-cycle factor. A cycle factor in a digraph $H$ is a spanning subdigraph of $H$ consisting of vertex-disjoint cycles. We will use the following characterization of Gutin [19, 20] of semicomplete multipartite digraph having a Hamilton path.
Theorem 1.5.
A semicomplete multipartite digraph has a Hamilton path if and only if it contains a 1-path-cycle factor. In polynomial time, one can decide whether a semicomplete multipartite digraph $D$ has a Hamilton path and find a Hamilton path in $D$, if it exists.
Let $D$ be a digraph. Its symmetric (0,1)-digraph is a digraph $\hat{D}$ obtained from $D$ by assigning cost 1 to the arcs of $D$ and adding arc $yx$ of cost 0 for every $xy\in A(D)$ such that $yx\not\in A(D)$. The cost of a subgraph $H$ of $\hat{D}$ is the sum of the costs of arcs of $H.$
The following result determines the maximum $\sigma_{\max}(P)$ for a Hamilton oriented path $P$ in a semicomplete multipartite digraph $D$ provided $D$ satisfies the HP-majority inequality. Theorem 1.6 generalizes Theorem 1.5, and is proved in Subsection 3.1.
Theorem 1.6.
Let $D$ be an $n$-vertex semicomplete multipartite digraph satisfying the HP-majority inequality and let $c^{pc}_{\max}$ be the maximum cost of a 1-path-cycle factor in $\hat{D}.$ Let $\sigma^{hp}_{\max}$ be the maximum $\sigma_{\max}(P)$ for a Hamilton oriented path $P$ in $D$. Then $\sigma^{hp}_{\max}=c^{pc}_{\max}$. Both $\sigma^{hp}_{\max}$ and a Hamilton oriented path $P$ with $\sigma_{\max}(P)=\sigma^{hp}_{\max}$ can be found in polynomial time.
While Bang-Jensen, Gutin and Yeo [8] proved that in polynomial time one can decide whether a semicomplete multipartite digraph has a Hamilton cycle, no characterezation of hamiltonian semicomplete multipartite digraphs has been obtained. The next theorem determines the maximum $\sigma_{\max}(C)$ for a Hamilton oriented cycle $C$ in a semicomplete multipartite digraph $D$ satisfying the HC-majority inequality.
Theorem 1.7.
Let $D$ be an $n$-vertex semicomplete multipartite digraph satisfying the HC-majority inequality and let $c^{cf}_{\max}$ be the maximum cost of a cycle factor in $\hat{D}.$ Let $\sigma^{hc}_{\max}$ be the maximum $\sigma_{\max}(C)$ for a Hamilton oriented cycle $C$ in $D$. Then $\sigma^{hc}_{\max}=c^{cf}_{\max}$ unless $c^{cf}_{\max}=n$ and $D$ is not hamiltonian, in which case $\sigma^{hc}_{\max}=n-1.$ Both $\sigma^{hc}_{\max}$ and a Hamilton oriented cycle $C$ with $\sigma_{\max}(C)=\sigma^{hc}_{\max}$ can be found in polynomial time.
Our proof of Theorem 1.7 uses the following theorem, which may be of independent interest.
Theorem 1.8.
Let $D$ be a semicomplete multipartite digraph and let ${F}$ be a 1-path-cycle factor in $D$. Let $P$ denote the path in ${F}$ and assume the initial and terminal vertices in $P$ belong to different partite sets. Then $D$ contains a Hamilton path whose initial and terminal vertices belong to different partite sets.
In turn, our proof of Theorem 1.8 is based on a structural result of Yeo [26].
For a digraph $D$ and $x\in V(D)$, let $N^{+}(x)=\{y\in V(D):\ xy\in A(D)\}$ and $N^{-}(x)=\{y\in V(D):\ yx\in A(D)\}.$ In Section 4 we consider locally semicomplete digraphs. A digraph $D$ is locally semicomplete if for every $x\in V(D)$, both $D[N^{+}(x)]$ and $D[N^{-}(x)]$ are semicomplete digraphs. A digraph $D$ is connected if $U(D)$ is connected. A digraph $D$ is strong if there is a path from $x$ to $y$ for every ordered pair $x,y$ of vertices of $D$. A strong component of a digraph $D$ is a maximal strong subgraph of $D$. It is well known [4, 5] that every connected locally semicomplete digraph has a Hamilton path. This implies that the problem of finding a Hamilton oriented path is trivial. Also, it is well known [4, 5] that every strong semicomplete digraph has a Hamilton cycle.
However, this does not imply a characterization of Hamilton oriented cycles with the maximum possible number of forward arcs. We provide such a characterization in Section 4. It is easy to see that we can order strong components $H_{1},\dots,H_{\ell}$ of a non-strongly connected digraph $H$ such that there is no arc from $H_{j}$ to $H_{i}$ for $i<j$. This is called an acyclic ordering of strong components of $H$. For a non-strongly connected locally semicomplete digraph, it is well known [4, 5] that an acyclic ordering of strong components is unique and has a few nice properties described in Section 4.
Let $S,T$ be vertex-disjoint subgraphs of a digraph $H$. A path $P=p_{1}\dots p_{t}$ of $H$ is an $(S,T)$-path if $p_{1}\in{V(S)}$, $p_{t}\in{V(T)}$ and all other vertices of $P$ are outside of $S\cup T$. Let $d(S,T)$ denote the length of a shortest $(S,T)$-path. Here is our characterization.
Theorem 1.9.
Let $D$ be a connected locally semicomplete digraph on $n\geq 3$ vertices. Let $C_{1},C_{2},\dots,C_{\ell}$ be the acyclic ordering of strong components of $D$. Let $\sigma^{hc}_{\max}$ be the maximum $\sigma_{\max}(R)$ for a Hamilton oriented cycle $R$ in $D$. The following now holds.
  • If $D$ is not strong and $U(D)$ is 2-connected, then $\sigma^{hc}_{\max}=n-d(C_{1},C_{\ell})$.
  • If $D$ is not strong and $U(D)$ is not 2-connected, then $D$ contains no oriented Hamilton cycle.
  • If $D$ is strong then $\sigma^{hc}_{\max}=n$.
Furthermore, in polynomial time, we can find a Hamilton oriented cycle of $D$ with the maximum number of forward arcs.
We conclude the paper in Section 5, where we discuss the complexity of finding a Hamilton oriented cycle with maximum number of forward arcs in other classes of digraphs.
2 Results in support of Conjectures 1.3 and 1.4
When we strengthen the degree sum condition in Conjecture 1.3, we deduce the following result which is the desired bound stated in the conjecture.
Theorem 2.1.
Let $D$ be an oriented graph on $n$ vertices with minimum degree $1<\delta<\frac{n}{2}$. If $d(u)+d(v)\geq n+\delta-2$ for each pair of non-adjacent vertices $u$ and $v$, then there exists a Hamilton oriented cycle $C$ in $D$ such that $\sigma_{\max}(C)\geq n-\delta$.
Observe that Theorem 2.1 implies that Conjecture 1.3 is true for $\delta=2$, since the condition in Theorem 2.1 is precisely the Ore-type condition.
The neighborhood of a vertex $v$ in a directed or undirected graph $D$, denoted by $N_{D}(v)$, is the set of vertices adjacent to $v$. And the closed neighborhood of $v$ denoted by $N_{D}[v]$, is defined as $N_{D}[v]=N_{D}(v)\cup\{v\}$. When the graph induced by the neighborhood of a vertex with the minimum degree in the graph is isomorphic to a tournament, we can derive a slightly weaker bound for Conjecture 1.3 based on this scenario.
Theorem 2.2.
Let $D$ be an oriented graph on $n$ vertices with minimum degree $3\leq\delta\leq\frac{n}{3}$, which satisfies the Ore-type condition. If there is a vertex $v_{0}\in V(D)$ such that $d(v_{0})=\delta$ and $D[N(v_{0})]$ is isomorphic to a tournament, then there is a Hamilton oriented cycle $C$ in $D$ such that $\sigma_{\max}(C)\geq n-\delta-2$.
By applying an approach from [17], we obtain the following approximate version of Conjecture 1.4.
Theorem 2.3.
For any integer $k\geq 0$ and oriented graph $D$ with order $n\geq 30+4(k-1)$. If $s^{*}(D)\geq 8k$, then there is a Hamilton oriented cycle $C$ in $D$ such that $\sigma_{\max}(C)\geq\lceil\frac{n+k}{2}\rceil$.
In the first subsection, we present the constructions that demonstrate the tightness of Conjectures 1.3 and 1.4. In the subsequent subsections, we provide proofs for the three theorems mentioned above.
2.1 Tightness of Conjectures 1.3 and 1.4
For an undirected graph $G$, let $\overline{G}$ denote its complement. The following construction shows that Conjecture 1.3 is sharp when $\delta<n/2$, for we cannot guarantee $D$ contains a Hamilton oriented cycle $C$ with $\sigma_{\max}(C)>n-\delta$.
Construction 2.4.
Given integers $n$ and $\delta$ with $\delta\geq 2$ and $n>2\delta$, let $G$ be a graph on $n$ vertices with vertex set $V(G)=A\cup B\cup\{v_{0}\}$, where $G[A]$ is isomorphic to $\overline{K}_{\delta}$, $G[B]$ is isomorphic to $K_{n-\delta-1}$, $N(v_{0})=A$, and every vertex in $A$ is adjacent to all vertices in $B$. Let $D$ be an orientation of $G$, such that every edge between $A$ and $B$ is oriented from $B$ to $A$, and every edge between $v_{0}$ and $A$ is oriented from $v_{0}$ to $A$ (see Figure 1).
It is not hard to check that the Ore-type condition holds for $D$ and $\delta(D)=\delta$. Let $C$ be a Hamilton oriented cycle in $D$. Observe that every vertex in $A$ is incident to two edges of $C$ which are oriented in opposite directions. Since $A$ is an independent set, then $\sigma_{\min}(C)\geq|A|=\delta$. It follows that $\sigma_{\max}(C)\leq n-\delta$.
[…]
Figure 1: Tight examples for Conjecture 1.3
The following construction shows the bound in Conjecture 1.4 is tight for all positive integer $n$ and integer $40\leq s^{*}(D)\leq n-$ 4 (note that if $s^{*}(D)\geq n-3$ then we are considering tournaments and the bounds are clearly tight for them).
Construction 2.5.
Given a positive integer $n$ and a non-negative integer $k$ with $n\geq k+4$, let $\epsilon\in\{0,1\}$ with $\epsilon\equiv n+k\pmod{2}$. Let $G$ be a graph on $n$ vertices with vertex set $V(G)=A\cup B$, where $G[A]$ is isomorphic to $\overline{K}_{\frac{n-k-\epsilon}{2}}$, $G[B]$ is isomorphic to $K_{\frac{n+k+\epsilon}{2}}$, and all possible edges with one end-vertex in $A$ and the other in $B$ exist, except $\epsilon$ edge. Let $D$ be an orientation of $G$, such that every edge between $A$ and $B$ is oriented from $B$ to $A$ (see Figure 2).
It is not hard to check that $s^{*}(D)=k$, and for every Hamilton oriented cycle $C$ in $D$, we have $\sigma_{\max}(C)\leq\lceil\frac{n+k}{2}\rceil$.
[…]
Figure 2: Tight examples for Conjecture 1.4
2.2 Proof of Theorem 2.1
Lemma 2.6.
Let $P=u_{1}u_{2}\ldots u_{n}$ be an oriented path and let $\{v_{1},v_{2},\dots,v_{k}\}$ be a set of vertices that are not in $V(P)$, satisfying $v_{i}$ adjacent to $u_{j}$ for each $i\in[k]$ and $j\in[n]$. If $n\geq k+2$, then there exists an oriented path $P^{\prime}$ which is generated by $P$, by adding all $v_{i}$ between $u_{a_{i}}$ and $u_{a_{i}+1}$ for $a_{i}\in[n-1]$ such that $a_{i}\neq a_{j}$ for $i\neq j$, satisfying $\sigma^{+}(P^{\prime})\geq\sigma^{+}(P)$. Such an oriented path $P^{\prime\prime}$ satisfying $\sigma^{-}(P^{\prime\prime})\geq\sigma^{-}(P)$ also exists.
Proof.
Note that it suffices to prove the existence of $P^{\prime}$ when $n=k+2$. For $k=1$, If the arc between $v_{1}$ and $u_{2}$ is $v_{1}u_{2}$, then $P^{\prime}=u_{1}v_{1}u_{2}u_{3}$ satisfies $\sigma^{+}(P^{\prime})\geq\sigma^{+}(P)$. Otherwise, the arc between $v_{1}$ and $u_{2}$ is $u_{2}v_{1}$, and then $P^{\prime}=u_{1}u_{2}v_{1}u_{3}$ satisfies $\sigma^{+}(P^{\prime})\geq\sigma^{+}(P)$. We now prove that for $k\geq 2$, the oriented path $P^{\prime}$ exists by induction on $k$.
Let $P_{k+2-t}=u_{1}u_{2}\ldots u_{t}$ be an oriented subpath of $P$, for $t\in[k+1]$. Note that $\sigma^{+}(P_{k+2-t})\geq\sigma^{+}(P)-(k+2-t)$. Suppose to the contrary that such $P^{\prime}$ does not exist for $k$. Then arcs $v_{i}u_{k+1}$ exist, for all $i\in[k]$. For otherwise, the arc between $v_{i_{0}}$ and $u_{k+1}$ is $u_{k+1}v_{i_{0}}$ for some $i_{0}\in[k]$. By induction, there is an oriented path $P^{\prime}_{1}$ which is generated by $P_{1}$, by adding $\{v_{1},v_{2},\dots,v_{k}\}\setminus\{v_{i_{0}}\}$ into $P_{1}$, satisfying $\sigma^{+}(P^{\prime}_{1})\geq\sigma^{+}(P_{1})$. Then set $P^{\prime}$ be the oriented path which consists of $P^{\prime}_{1}$, the arc $u_{k+1}v_{i_{0}}$ and the arc between $v_{i_{0}}$ and $u_{k+2}$. Then $\sigma^{+}(P^{\prime})\geq\sigma^{+}(P^{\prime}_{1})+1\geq\sigma^{+}(P)$, which is a contradiction. Then arcs $v_{i}u_{k}$ exist, for all $i\in[k]$. For otherwise, the arc between $v_{i_{0}}$ and $u_{k}$ is $u_{k}v_{i_{0}}$ for some $i_{0}\in[k]$. Then we arbitrarily choose $v_{i_{1}}$ for some $i_{1}\in[k]$ and $i_{1}\neq i_{0}$. By induction, there is an oriented path $P^{\prime}_{2}$ which is generated by $P_{2}$, by adding $\{v_{1},v_{2},\dots,v_{k}\}\setminus\{v_{i_{0}},v_{i_{1}}\}$ into $P_{2}$, satisfying $\sigma^{+}(P^{\prime}_{2})\geq\sigma^{+}(P_{2})$. Then set $P^{\prime}$ be the oriented path which consists of $P^{\prime}_{2}$, the directed path $u_{k}v_{i_{0}}u_{k+1}$, the arc $v_{i_{1}}u_{k+1}$ and the arc between $v_{i_{1}}$ and $u_{k+2}$. Then $\sigma^{+}(P^{\prime})\geq\sigma^{+}(P^{\prime}_{2})+2\geq\sigma^{+}(P)$, which is a contradiction. Then by the similar discussion, arcs $v_{i}u_{j}$ exists, for all $i\in[k]$ and $j\in[k+1]\setminus\{1\}$. Set $P^{\prime}$ be the oriented path which is generated by $P$, by adding all $v_{i}$ between $u_{i}$ and $u_{i+1}$ for $i\in[k]$. Then $\sigma^{+}(P^{\prime})\geq\sigma^{+}(P)$, which is a contradiction. ∎
Lemma 2.6 allows us to add a sufficient number of vertices to an oriented path or cycle, but it does not decrease the discrepancy of the oriented path or cycle. With Lemma 2.6 established, we are now ready to prove Theorem 2.1.
Theorem 2.1. Let $D$ be an oriented graph on $n$ vertices with minimum degree $1<\delta<\frac{n}{2}$. If $d(u)+d(v)\geq n+\delta-2$ for each pair of non-adjacent vertices $u$ and $v$, then there exists a Hamilton oriented cycle $C$ in $D$ such that $\sigma_{\max}(C)\geq n-\delta$.
Proof.
Let $v_{0}\in V(D)$ and $d(v_{0})=\delta$. Set $N(v_{0})=\{v_{1},v_{2},\ldots,v_{\delta}\}$ and $U=V(D)\setminus N[v_{0}]=\{u_{1},u_{2},\ldots,u_{n-\delta-1}\}$. Since $d(u)+d(v)\geq n+\delta-2$ for any non-adjacent $u,v\in V(D)$, then $d(u)\geq n-2$ for all $u\in U$, which means $u$ is adjacent to all vertices except $v_{0}$ in $D$.
For $n=5$, there is a Hamilton oriented cycle $C$ in $D$ such that $\sigma_{\max}(C)\geq 3\geq n-\delta$. For $n\geq 6$, since $D[U]$ is isomorphic to a tournament of order $n-\delta-1$, Theorem 1.2 implies that there is a Hamilton oriented cycle $C_{0}$ in $D[U]$ such that $\sigma_{\max}(C_{0})\geq n-\delta-2$. Without loss of generality, we assume that $C_{0}=u_{1}u_{2}\ldots u_{n-\delta-1}u_{1}$, $\sigma_{\max}(C_{0})=\sigma^{+}(C_{0})$ and $u_{2}u_{1}$ is the only backward arc in $C_{0}$ (If $\sigma_{\max}(C_{0})=n-\delta-2$). For each $i\in[n-\delta-1]$, set Hamilton oriented cycles $C^{1}_{i,i+1}=u_{i}v_{1}v_{0}v_{2}u_{i+1}u_{i+2}\dots u_{i-1}u_{i}\pmod{n-\delta-1}$ and $C^{2}_{i,i+1}=u_{i}v_{2}v_{0}v_{1}u_{i+1}u_{i+2}\dots u_{i-1}u_{i}\pmod{n-\delta-1}$ in $D[U\cup\{v_{0},v_{1},v_{2}\}]$.
Claim 1.
There exists $C^{t}_{i,i+1}$ such that $\sigma^{+}(C^{t}_{i,i+1})\geq n-\delta$ for some $i\in[n-\delta-1]$ and $t\in[2]$.
Suppose that $\sigma^{+}(C^{t}_{i,i+1})\leq n-\delta-1$ for each $i\in[n-\delta-1]$ and $t\in[2]$. We may assume $\{v_{1}v_{0},v_{2}v_{0}\}\subset A(D)$ or $\{v_{0}v_{1},v_{0}v_{2}\}\subset A(D)$, since otherwise $\sigma^{+}(C^{1}_{1,2})\geq n-\delta$ if $v_{1}v_{0}v_{2}$ is a directed path or $\sigma^{+}(C^{2}_{1,2})\geq n-\delta$ if $v_{2}v_{0}v_{1}$ is a directed path. Without loss of generality, assume $\{v_{1}v_{0},v_{2}v_{0}\}\subset A(D)$. Since $v_{1}v_{0}\in A(D)$, then $\sigma^{+}(C^{1}_{1,2})\geq n-\delta-1$ and then $\sigma^{+}(C^{1}_{1,2})=n-\delta-1$. Thus $\{v_{1}u_{1},u_{2}v_{2}\}\subset A(D)$. Similarly, since $v_{2}v_{0}\in A(D)$, then $\sigma^{+}(C^{2}_{1,2})=n-\delta-1$ and $\{v_{2}u_{1},u_{2}v_{1}\}\subset A(D)$. Furthermore, since $\{u_{2}v_{1},v_{1}v_{0}\}\subset A(D)$, then $\sigma^{+}(C^{1}_{2,3})\geq n-\delta-1$. Thus $u_{3}v_{2}\in A(D)$. And since $\{u_{2}v_{2},v_{2}v_{0}\}\subset A(D)$, then $\sigma^{+}(C^{2}_{2,3})\geq n-\delta-1$ and $u_{3}v_{1}\in A(D)$. By discussing $C^{t}_{i,i+1}$ for every $i\in[n-\delta-1]$ and $t\in[2]$ in a similar way, we have $\{u_{i+1}v_{1},u_{i+1}v_{2}\}\subset A(D)$ for every $i\in[n-\delta-1]$. Then $\{u_{1}v_{1},u_{1}v_{2},v_{1}u_{1},v_{2}u_{1}\}\subset A(D)$, which contradicts the condition that $D$ is an oriented graph. Thus Claim 1 holds.
Let $C^{t_{0}}_{i_{0},i_{0}+1}$ be the oriented cycle such that $\sigma^{+}(C^{t_{0}}_{i_{0},i_{0}+1})\geq n-\delta$ as in Claim 1. Set the oriented path $P=u_{i_{0}+1}u_{i_{0}+2}\dots u_{i_{0}}\pmod{n-\delta-1}$. Since $\delta-2\leq(n-\delta-1)-2$, Lemma 2.6 implies that there exists an oriented path $P^{\prime}$ with $V(P^{\prime})=V(P)\cup\{v_{3},\dots,v_{\delta}\}$ and the same initial and terminal vertices as $P$, satisfying $\sigma^{+}(P^{\prime})\geq\sigma^{+}(P)$. Let $C$ be the Hamilton oriented cycle in $D$, whose arc set consists of $A(C^{t_{0}}_{i_{0},i_{0}+1})\setminus A(P)$ and $A(P^{\prime})$. Then $\sigma_{\max}(C)\geq\sigma^{+}(C^{t_{0}}_{i_{0},i_{0}+1})=n-\delta$. ∎
2.3 Proof of Theorem 2.2
Firstly, we need the following lemma:
Lemma 2.7.
Let $D$ be a tournament on $n\geq 3$ vertices. For a Hamilton oriented cycle $C=v_{1}v_{2}\dots v_{i-1}v_{i}v_{i+1}\dots v_{j-1}v_{j}v_{j+1}\dots v_{n}v_{1}$ in $D$, let $C^{\prime}$ be a Hamilton oriented cycle which is generated by $C$, satisfying $C^{\prime}=v_{1}v_{2}\dots v_{i-1}v_{j}v_{i+1}\dots v_{j-1}v_{i}v_{j+1}\dots\\ v_{n}v_{1}$. Then $\sigma^{+}(C^{\prime})\geq\sigma^{+}(C)-4$ and $\sigma^{-}(C^{\prime})\geq\sigma^{-}(C)-4$. In particular, if $v_{i}$ and $v_{j}$ are adjacent in $C$, then $\sigma^{+}(C^{\prime})\geq\sigma^{+}(C)-3$ and $\sigma^{-}(C^{\prime})\geq\sigma^{-}(C)-3$.
Proof.
$C^{\prime}$ is definitely an oriented cycle, as $D$ is a tournament. Since all vertices except $v_{i}$ and $v_{j}$ appear in the same order in $C$ and $C^{\prime}$ and there are at most four arcs in $C$ incident with $v_{i}$ or $v_{j}$, then $\sigma^{+}(C^{\prime})\geq\sigma^{+}(C)-4$ and $\sigma^{-}(C^{\prime})\geq\sigma^{-}(C)-4$. ∎
Theorem 2.2. Let $D$ be an oriented graph on $n$ vertices with minimum degree $3\leq\delta\leq\frac{n}{3}$, which satisfies the Ore-type condition. If there is a vertex $v_{0}\in V(D)$ such that $d(v_{0})=\delta$ and $D[N(v_{0})]$ is isomorphic to a tournament, then there is a Hamilton oriented cycle $C$ in $D$ such that $\sigma_{\max}(C)\geq n-\delta-2$.
Proof.
Suppose that $v_{0}\in V(D)$ with $d(v_{0})=\delta$ and $D[N(v_{0})]$ is isomorphic to a tournament. We denote $V(D)\backslash N[v_{0}]$ by $W$. Then $|W|=n-\delta-1$. Since for every vertex $w\in W$, we have $d(w)\geq n-\delta$ and $n-2\delta\leq d_{D[W]}(w)\leq n-\delta-2$, then $w$ must be adjacent to at least two vertices in $N(v_{0})$.
Since $\delta(D[W])\geq n-2\delta\geq\frac{n-\delta-1}{2}=\frac{|W|}{2}$ when $\delta\leq\frac{n}{3}$, Theorem 1.2 implies that there is a Hamilton oriented cycle $C_{1}$ in $D[W]$ such that $\sigma_{\max}(C_{1})\geq\delta(D[W])$. Set $C_{1}=w_{1}w_{2}\ldots w_{n-\delta-1}w_{1}$. Similarly, there exists an oriented Hamilton cycle $C_{2}=v_{0}v_{1}v_{2}\ldots v_{\delta}v_{0}$ with $\sigma_{\max}(C_{2})\geq\delta$ in $D[N[v_{0}]]$. Set $w_{k}\in W$ with $d_{D[W]}(w_{k})=\delta(D[W])$. Since $V(C_{1})\cap V(C_{2})=\emptyset$, without loss of generality, we can set $\sigma^{+}(C_{i})=\sigma_{\max}(C_{i})$, for each $i\in[2]$.
Claim 2.
If there exist $v_{j}$ and $v_{j+1}$ for $j\in[\delta-1]$ in $C_{2}$ that are adjacent to $w_{k+1}\pmod{n-\delta-1}$ and $w_{k}$ respectively, then there is a Hamilton oriented cycle $C$ in $D$ such that $\sigma_{\max}(C)\geq n-\delta-2$.
Set $C=w_{1}\ldots w_{k}v_{j+1}v_{j+2}\ldots v_{\delta}v_{0}v_{1}\dots v_{j}w_{k+1}\ldots w_{n-\delta-1}w_{1}$, which is generated by $C_{1}$ by deleting the arc between $w_{k}$ and $w_{k+1}$, and adding the arc between $w_{k}$ and $v_{j+1}$, the arc between $w_{k+1}$ and $v_{j}$, and the longer part of $C_{2}$ from $v_{j+1}$ to $v_{j}$. Then $\sigma_{\max}(C)\geq\sigma^{+}(C)\geq n-2\delta-1+\delta-1=n-\delta-2$. Thus Claim 2 holds.
Then we may assume that for each $j\in[\delta-1]$, if $w_{k+1}$ is adjacent to $v_{j}$, then $w_{k}$ is not adjacent to $v_{j+1}$ in the following discussion. Then $w_{k}$ is not adjacent to at least one vertex in $V(C_{2})\setminus\{v_{0}\}$ since $w_{k+1}$ has at least two neighbors in $V(C_{2})\setminus\{v_{0}\}$. In this case, $\delta(D[W])\geq n-2\delta+1$.
Claim 3.
If $v_{1}$ and $v_{\delta}$ are adjacent to $w_{k}$ and $w_{k+1}\pmod{n-\delta-1}$ respectively, then there is a Hamilton oriented cycle $C$ in $D$ such that $\sigma_{\max}(C)\geq n-\delta-2$.
Set $C^{\prime}=w_{1}\ldots w_{k}v_{1}v_{2}\ldots v_{\delta}w_{k+1}\ldots w_{n-\delta-1}w_{1}$, which is generated by $C_{1}$ by deleting the arc between $w_{k}$ and $w_{k+1}$, and adding the arc between $w_{k}$ and $v_{1}$, the arc between $w_{k+1}$ and $v_{\delta}$, and the part of $C_{2}$ from $v_{1}$ to $v_{\delta}$. Then $\sigma^{+}(C^{\prime})\geq n-2\delta+1-1+\delta-2=n-\delta-2$. By Lemma 2.6, we can add $v_{0}$ to $C^{\prime}$ when $\delta\geq 3$, to obtain the desired Hamilton oriented cycle $C$ in $D$ such that $\sigma_{\max}(C)\geq\sigma_{\max}(C^{\prime})\geq n-\delta-2$. Thus Claim 3 holds.
When $\delta(D[W])=n-2\delta+1$, then $w_{k}$ is not adjacent to at most one vertex in $V(C_{2})\setminus\{v_{0}\}$. By the assumption before, $w_{k}$ is not adjacent to exactly one vertex $v_{i_{1}}$ in $V(C_{2})\setminus\{v_{0}\}$. Since $w_{k+1}$ is adjacent to at least two vertices in $V(C_{2})\setminus\{v_{0}\}$, let $v_{j_{1}}$ and $v_{j_{2}}$ be adjacent to $w_{k+1}$, in which $j_{1}<j_{2}$. By the assumption before, we have $i_{1}=j_{1}+1$ and $j_{2}=\delta$. Then $v_{1}$ and $v_{\delta}$ are adjacent to $w_{k}$ and $w_{k+1}$ respectively. By Claim 3, we are done.
We may also assume that either $w_{k}$ is not adjacent to $v_{1}$ or $w_{k+1}$ is not adjacent to $v_{\delta}$ in the following discussion. In this case, $\delta(D[W])\geq n-2\delta+2$.
Claim 4.
If there exist $v_{j}$ and $v_{j+1}$ for $j\in[\delta-1]$ in $C_{2}$ that are adjacent to $w_{k}$ and $w_{k+1}\pmod{n-\delta-1}$ respectively, then there is a Hamilton oriented cycle $C$ in $D$ such that $\sigma_{\max}(C)\geq n-\delta-2$.
Let $C^{\prime}_{2}$ be an oriented cycle which is generated by $C_{2}$ in Lemma 2.7, by changing $v_{j}$ and $v_{j+1}$ in $C_{2}$. Then by Lemma 2.7, $\sigma^{+}(C_{2}^{\prime})\geq\sigma^{+}(C_{2})-3\geq\delta-3$. Set $C=w_{1}\ldots w_{k}v_{j}v_{j+2}\ldots v_{\delta}v_{0}v_{1}\dots v_{j-1}v_{j+1}w_{k+1}\ldots w_{n-\delta-1}w_{1}$, which is generated by $C_{1}$ by deleting the arc between $w_{k}$ and $w_{k+1}$, and adding the arc between $w_{k}$ and $v_{j}$, the arc between $w_{k+1}$ and $v_{j+1}$, and the longer part of $C^{\prime}_{2}$ from $v_{j}$ to $v_{j+1}$. Then $\sigma_{\max}(C)\geq\sigma^{+}(C)\geq n-2\delta+2-1+\delta-3=n-\delta-2$. Thus Claim 4 holds.
Then based on the former two assumptions, we may assume that for each $j\in[\delta-1]$, if $w_{k+1}$ is adjacent to $v_{j+1}$, then $w_{k}$ is not adjacent to $v_{j}$ in the following discussion.
When $\delta(D[W])=n-2\delta+2$, then $w_{k}$ is not adjacent to at most two vertices in $V(C_{2})\setminus\{v_{0}\}$. By the assumptions before, $w_{k}$ is not adjacent to exactly two vertices $v_{i_{1}}$ and $v_{i_{2}}$ in $V(C_{2})\setminus\{v_{0}\}$, in which $i_{1}<i_{2}$. Since $w_{k+1}$ is adjacent to at least two vertices in $V(C_{2})\setminus\{v_{0}\}$, let $v_{j_{1}}$ and $v_{j_{2}}$ be adjacent to $w_{k+1}$, in which $j_{1}<j_{2}$. By the assumptions before, $i_{1}=j_{1}+1=j_{2}-1$ and $i_{2}=j_{2}+1$ and $j_{1}=1$, or $i_{1}=j_{1}-1$ and $i_{2}=j_{1}+1=j_{2}-1$ and $j_{2}=\delta$. Then both of $w_{k}$ and $w_{k+1}$ are adjacent to $v_{1}$ and $v_{3}$, or $v_{\delta-2}$ and $v_{\delta}$. Let $C^{\prime}_{2}$ be an oriented cycle which is generated by $C_{2}$ in Lemma 2.7, by changing $v_{2}$ and $v_{3}$ in $C_{2}$, or by changing $v_{\delta-2}$ and $v_{\delta-1}$ in $C_{2}$. Then by Lemma 2.7, $\sigma^{+}(C_{2}^{\prime})\geq\sigma^{+}(C_{2})-3\geq\delta-3$. Set $C=w_{1}\ldots w_{k}v_{3}v_{2}v_{4}\ldots v_{\delta}v_{0}v_{1}w_{k+1}\ldots w_{n-\delta-1}w_{1}$, which is generated by $C_{1}$ by deleting the arc between $w_{k}$ and $w_{k+1}$, and adding the arc between $w_{k}$ and $v_{3}$, the arc between $w_{k+1}$ and $v_{1}$, and the longer part of $C^{\prime}_{2}$ from $v_{3}$ to $v_{1}$, or set $C=w_{1}\ldots w_{k}v_{\delta}v_{0}v_{1}\ldots v_{\delta-3}v_{\delta-1}v_{\delta-2}w_{k+1}\ldots w_{n-\delta-1}w_{1}$, which is generated by $C_{1}$ by deleting the arc between $w_{k}$ and $w_{k+1}$, and adding the arc between $w_{k}$ and $v_{\delta}$, the arc between $w_{k+1}$ and $v_{\delta-2}$, and the longer part of $C^{\prime}_{2}$ from $v_{\delta}$ to $v_{\delta-2}$. Then $\sigma_{\max}(C)\geq\sigma^{+}(C)\geq n-2\delta+2-1+\delta-3=n-\delta-2$.
When $\delta(D[W])\geq n-2\delta+3$, assume that the vertices in $V(C_{2})$ which are adjacent to $w_{k+1}$ are $\{v_{j_{1}},\ldots,v_{j_{a}}\}$, in which $a\geq 2$ and $j_{1}<j_{2}<\dots<j_{a}$. Then $w_{k}$ is not adjacent to $v_{j_{1}+1}$ by the assumptions before. Set $v_{i_{1}}\in V(C_{2})\setminus\{v_{0}\}$ that is adjacent to $w_{k}$. Without loss of generality, assume $i_{1}<j_{1}-1$. Let $C^{\prime}_{2}$ be an oriented cycle which is generated by $C_{2}$ in Lemma 2.7, by changing $v_{i_{1}}$ and $v_{j_{1}+1}$ in $C_{2}$. Then by Lemma 2.7, $\sigma^{+}(C_{2}^{\prime})\geq\delta-4$. Then set $C^{\prime}=w_{1}\ldots w_{k}v_{i_{1}}v_{j_{1}+2}v_{j_{1}+3}\dots v_{\delta}v_{0}v_{1}\dots\\ v_{i_{1}-1}v_{j_{1}+1}v_{i_{1}+1}\ldots v_{j_{1}}w_{k+1}\ldots w_{n-\delta-1}w_{1}$, which is generated by $C_{1}$ by deleting the arc between $w_{k}$ and $w_{k+1}$, and adding the arc between $w_{k}$ and $v_{i_{1}}$, the arc between $w_{k+1}$ and $v_{j_{1}}$, and the longer part of $C^{\prime}_{2}$ from $v_{i_{1}}$ to $v_{j_{1}}$. Then $\sigma_{\max}(C)\geq\sigma^{+}(C)\geq n-2\delta+3-1+\delta-4=n-\delta-2$. ∎
2.4 Proof of Theorem 2.3
The following lemma informs us of a lower bound on the size $a(D)$ of the arc set of $D$. For an undirected graph $G$, let $e(G)=|E(G)|.$
Lemma 2.8.
Let $D$ be an oriented graph and $t$ a non-negative integer. If $s^{*}(D)\geq t$, then $a(D)\geq\frac{n(n+t)}{4}$.
Proof.
As $D$ is an oriented graph, we only need to show that the underlying graph $G$ of $D$ has at least $\frac{n(n+t)}{4}$ edges. Consider the complement $\overline{G}$ of $G$. On the one hand, since $s^{*}(D)\geq t$, then
$\sum_{uv\in E(\overline{G})}(d_{\overline{G}}(u)+d_{\overline{G}}(v))$ $=$ $\sum_{uv\not\in E(G)}(2n-2-(d_{G}(u)+d_{G}(v)))$ (1)
$\leq$ $e(\overline{G})(n-2-t).$
On the other hand, by Cauchy-Schwarz inequality, we have that
$\sum_{uv\in E(\overline{G})}(d_{\overline{G}}(u)+d_{\overline{G}}(v))=\sum_{u\in V(\overline{G})}d_{\overline{G}}^{2}(u)\geq\frac{1}{n}\left(\sum_{u\in V(\overline{G})}d_{\overline{G}}(u)\right)^{2}=\frac{4e^{2}(\overline{G})}{n}.$ (2)
By (1) and (2), we have $e(\overline{G})\leq n(n-2-t)/4,$ and therefore $e(G)={n\choose 2}-e(\overline{G})\geq\frac{n(n+t)}{4}$, which completes the proof. ∎
We call an undirected graph $G$ a diamond, if $G$ is isomorphic to $K_{4}^{-}$, where $K_{4}^{-}$ is obtained from $K_{4}$ by removing one edge. Let $V(G)=\{a,b,c,d\}$, $d(a)=d(b)=3$ and $d(c)=d(d)=2$. Given an orientation of $G$, it is a good diamond if $c$ and $d$ relate in the same way to $a,b$ (see Figure 3). Let $D$ be a good diamond. Note that there are oriented paths $P_{cd}$ and $P_{dc}$ of length 3 from $c$ to $d$ and from $d$ to $c$ respectively in $D$, such that $\sigma^{+}(P_{cd})\geq 2$ and $\sigma^{+}(P_{dc})\geq 2$.
The following lemmas can be found in [17]. While the second lemma is a slight generalization of the result by Pósa [25], it can be shown by slightly modifying his proof for graphs $G$ with $\delta(G)\geq(n+k)/2$. However, for the sake of completion, we provide a proof here.
Lemma 2.9 ([17]).
Let $k\geq 1$, and let $G$ be a graph with $|V(G)|=n\geq 30+4(k-1)$ and $e(G)\geq\frac{n^{2}}{4}+2(k-1)n-4k^{2}+6k-1$. Then every orientation of $G$ contains $k$ vertex-disjoint good diamonds.
$a$$b$$c$$d$
$a$$b$$c$$d$
$a$$b$$c$$d$
$a$$b$$c$$d$
Figure 3: The four good diamonds
Lemma 2.10.
Let $t\geq 0$ and let $G$ be a graph with $n$ vertices and $s^{*}(G)\geq t$. Let $E\subseteq E(G)$ be the edge set of a path forest of size at most $t$. Then there exists a Hamilton cycle in G which uses all edges in $E$.
Proof.
Suppose $G$ is a maximal counterexample to this lemma, which means by adding any new edge $uv$, the lemma holds for $G+uv$. Note that $G$ cannot be a complete graph as otherwise this lemma would have held for $G$. Let $v_{1}$ and $v_{n}$ be a pair of non-adjacent vertices. As the lemma holds for $G+v_{1}v_{n}$ and not for $G$, there is a Hamilton cycle $C=v_{1}v_{2}\dots v_{n}v_{1}$ where $E\subseteq E(C)$. Let $P$ be the Hamilton path $P=C-\{v_{1}v_{n}\}$ in $G$.
For a property $\cal P$, let $I({\cal P})$ be the indicator of $\cal P$ i.e. $I({\cal P})=1$ if $\cal P$ holds and $I({\cal P})=0$, otherwise. For every $i\in[n]$, let $I(v_{1}v_{i})$ and $I(v_{n}v_{i})$ be the indicators for whether $v_{1}v_{i}\in E(G)$ and $v_{n}v_{i}\in E(G)$ respectively. Let $S=E(P)\setminus E$. Thus, $|S|=n-1-|E|$. Note that for every $v_{i-1}v_{i}\in S$, we have that
$I(v_{1}v_{i})+I(v_{n}v_{i-1})\leq 1,$ (3)
as it is obvious if $i=2$ or $n$ (since $I(v_{1}v_{n})=0$) and if $i\in\{3,4,\dots,n-1\}$ and $I(v_{1}v_{i})+I(v_{n}v_{i-1})=2$, then $v_{1}v_{i}v_{i+1}\dots v_{n}v_{i-1}v_{i-2}\dots v_{1}$ is a Hamilton cycle containing all edges in $E$, a contradiction. Thus, on the one hand, by (3), we have that
$\sum_{v_{i-1}v_{i}\in E(P)}(I(v_{1}v_{i})+I(v_{n}v_{i-1}))\leq|S|+2|E|=n+|E|-1\leq n+t-1.$ (4)
On the other hand,
$\sum_{v_{i-1}v_{i}\in E(P)}(I(v_{1}v_{i})+I(v_{n}v_{i-1}))$ $=$ $\sum_{i=1}^{n-1}I(v_{n}v_{i})+\sum_{i=2}^{n}I(v_{1}v_{i})$
$=$ $d(v_{n})+d(v_{1})\geq n+t,$
which contradicts (4). This completes the proof. ∎
Now we give the proof of Theorem 2.3.
Theorem 2.3. For any integer $k\geq 0$ and oriented graph $D$ with order $n\geq 30+4(k-1)$. If $s^{*}(D)\geq 8k$, then there is a Hamilton oriented cycle $C$ in $D$ such that $\sigma_{\max}(C)\geq\lceil\frac{n+k}{2}\rceil$.
Proof.
Let $D$ be an oriented graph with order $n\geq 30+4(k-1)$ and $s^{*}(D)\geq 8k$. Then by Lemma 2.8, we have
$a(D)\geq\frac{n(n+8k)}{4}\geq\frac{n^{2}}{4}+2(k-1)n-4k^{2}+6k-1.$
Thus, by Lemma 2.9, $D$ contains $k$ vertex-disjoint good diamonds. Let $c_{i}$ and $d_{i}$ be the two vertices with degree $2$ on diamond $i$, and $P_{i}$ and $Q_{i}$ oriented paths of length 3 from $c_{i}$ to $d_{i}$ and $d_{i}$ to $c_{i}$ respectively, satisfying $\sigma^{+}(P_{i})\geq 2$ and $\sigma^{+}(Q_{i})\geq 2$. By Lemma 2.10, there is a Hamilton oriented cycle $C$ containing all arcs in $P_{i}$ (for all $i\in[k]$). Without loss of generality, we assume that most of the arcs in $A(C)\setminus(\cup_{i=1}^{k}A(P_{i}))$ go clockwise. Then, we can replace some $P_{i}$ by $Q_{i}$ with most of the arcs going clockwise if necessary. Thus, there are at least $\frac{n-3k}{2}+2k\geq\frac{n+k}{2}$ arcs going clockwise and therefore $\sigma_{\max}(C)\geq\frac{n+k}{2}$. ∎
3 Hamilton oriented paths and cycles with maximum number of forward arcs in semicomplete multipartite digraphs
The following result gives a characterization of semicomplete multipartite digraphs with a Hamilton oriented cycle (path, respectively).
Proposition 3.1.
A semicomplete multipartite digraph $D$ with partite sets of sizes $n_{1},n_{2},\dots,n_{p}$ has a Hamilton oriented cycle (path, respectively) if and only if $D$ satisfies the HC-majority inequality (the HP-majority inequality, respectively).
Proof.
Observe that $D$ has a Hamilton oriented cycle if and only if $K_{n_{1},n_{2},\dots,n_{p}}$ has a Hamilton cycle. Note that if $K_{n_{1},n_{2},\dots,n_{p}}$ does not satisfy the HC-majority inequality, then it has no Hamilton cycle. If $K_{n_{1},n_{2},\dots,n_{p}}$ satisfies the HC-majority inequality, then it has a Hamilton cycle by Dirac’s theorem [11].
Observe that $D$ has a Hamilton oriented path if and only if the digraph $D^{\prime}$ obtained from $D$ by adding a new vertex $x$ and all arcs of the form $xv,vx$ for $v\in V(D)$ has a Hamilton oriented cycle which is if and only if $K_{n_{1},n_{2},\dots,n_{p},1}$ has a Hamilton cycle. Thus, $D$ has a Hamilton oriented path if and only if it satisfies the HP inequality. ∎
3.1 Hamilton Oriented Paths
Let $D$ be a digraph. Recall that its symmetric (0,1)-digraph is a digraph $\hat{D}$ obtained from $D$ by assigning cost 1 to the arcs of $D$ and adding arc $yx$ of cost 0 for every $xy\in A(D)$ such that $yx\not\in A(D)$. The cost of a subgraph $H$ of $\hat{D}$ is the sum of the costs of arcs of $H.$ The following theorem characterizes the discrepancy of Hamilton oriented paths in semicomplete multipartite digraphs by the cost of path-cycle factors in its symmetric (0,1)-digraph.
Theorem 1.6. Let $D$ be a semicomplete multipartite digraph on $n$ vertices, satisfying the HP-majority inequality and let $c^{pc}_{\max}$ be the maximum cost of a 1-path-cycle factor in $\hat{D}.$ Let $\sigma^{hp}_{\max}$ be the maximum $\sigma_{\max}(P)$ for a Hamilton oriented path $P$ in $D$. Then $\sigma^{hp}_{\max}=c^{pc}_{\max}$. Both $\sigma^{hp}_{\max}$ and a Hamilton oriented path $P$ with $\sigma_{\max}(P)=\sigma^{hp}_{\max}$ can be found in polynomial time.
Proof.
Note that a Hamilton path $P=x_{1}x_{2}\dots x_{n}$ in $\hat{D}$ can be transformed into a Hamilton oriented path of $D$ by replacing every arc $x_{i}x_{i+1}$ of zero-cost by $x_{i+1}x_{i}$. Also by the construction of $\hat{D}$, $\sigma^{hp}_{\max}$ is equal to the maximum cost of a Hamilton path in $\hat{D}.$
Observe that a Hamilton path of $\hat{D}$ is a 1-path-cycle factor of $\hat{D}$. Thus, the maximum cost of a Hamilton path of $\hat{D}$ is smaller or equal to $c^{pc}_{\max}.$ Let $F$ be a 1-path-cycle factor of $\hat{D}$ of maximum cost and let $D_{F}$ be a spanning subdigraph of $\hat{D}$ with $A(D_{F})=A(D)\cup A(F).$ By Theorem 1.5, $D_{F}$ has a Hamilton path $P$. Note that the number of zero-cost arcs in $P$ cannot be larger than that in $F$, which means that the cost of $P$ is not smaller than that of $F$. Thus, the maximum cost of a Hamilton path of $\hat{D}$ is equal to $c^{pc}_{\max}$.
By Theorem 1.5 and the proof above, given a maximum cost 1-path-cycle factor in $\hat{D},$ we can find both $\sigma^{hp}_{\max}$ and a Hamilton oriented path $P$ with $\sigma_{\max}(P)=\sigma^{hp}_{\max}$ in polynomial time. To complete the proof of this theorem, we will describe how one can find a maximum cost 1-path-cycle factor in $\hat{D}$ in polynomial time. Let $\hat{D}^{\prime}$ be obtained from $\hat{D}$ by exchanging costs: 0 to 1 and 1 to 0 simultaneously. Observe that a maximum cost 1-path-cycle factor in $\hat{D}$ is a minimum cost 1-path-cycle factor in $\hat{D}^{\prime}$ and vice versa. By a remark in the last paragraph of Sec. 4.11.3 of [6] using minimum cost flows, one can find a minimum cost 1-path-cycle factor in an arc-weighted digraph in polynomial time. Since details on how to do it are omitted in [6], we give them below.
Add new vertices $s$ and $t$ to $\hat{D}^{\prime}$ together with arcs $sv$ and $vt$ of cost 0 for all vertices $v\in V(D)$ and assign lower and upper bound 1 to every vertex in $V(D)\cup\{s,t\}.$ In this network $N$ in polynomial time, we can find a minimum cost $(s,t)$-flow $f_{\min}$ of value 1. Note that all arcs of $N-\{s,t\}$ with flow of value 1 form a minimum cost 1-path-cycle factor in $\hat{D}^{\prime}$ which is a maximum cost 1-path-cycle factor in $\hat{D}.$
3.2 Hamilton Oriented Cycles
Recall the following theorem which is important for the main result of this section.
Theorem 3.2 ([8, 27]).
There is a polynomial-time algorithm for deciding whether a semicomplete multipartite digraph has a Hamilton cycle.
If a vertex $v$ belongs to a cycle $C$ then we denote the successor of $v$ on the cycle by $v_{C}^{+}$ and the predecessor by $v_{C}^{-}$. When $C$ is clear from the context, we may omit the subscript $C$. Let $C_{1}$ and $C_{2}$ be two disjoint cycles in a semicomplete multipartite digraph $D$. Suppose that $D$ has some partite set $V_{i}$, such that the following holds:
For every arc $u_{2}v_{1}$ from $C_{2}$ to $C_{1}$ we have $\{u_{2}^{+},v_{1}^{-}\}\subseteq V_{i}$, where $u_{2}^{+}$ is the successor of $u_{2}$ on $C_{2}$ and $v_{1}^{-}$ is the predecessor of $v_{1}$ on $C_{1}$.
Then we say that $C_{1}$ $V_{i}$-weakly-dominates $C_{2}$ and denote this by $C_{1}\rightsquigarrow_{V_{i}}C_{2}$. If $C_{1}\rightsquigarrow_{V_{i}}C_{2}$ for some $i$ then we also say that $C_{1}$ weakly-dominates $C_{2}$ and denote this simply by $C_{1}\rightsquigarrow C_{2}$.
See Figure 4 for an illustration of this definition. For example, in Figure 4, $w_{1}y_{2}$ is the only arc from $C_{3}$ to $C_{2}$ and $\{w_{1}^{+},y_{2}^{-}\}\in V_{3}$ (as $w_{1}^{+}=w_{2}$ and $y_{2}^{-}=y_{1}$). Therefore, $C_{2}\rightsquigarrow_{V_{3}}C_{3}$.
[…]
Figure 4: Arcs from cycle $C_{i}$ to $C_{j}$, for $1\leq i<j\leq 3$ are not shown. Note that $C_{1}\rightsquigarrow_{V_{1}}C_{2}$ and $C_{2}\rightsquigarrow_{V_{3}}C_{3}$. As there are no arcs from $C_{3}$ to $C_{1}$, we have $C_{1}\rightsquigarrow_{V_{i}}C_{3}$ for all $i=1,2,3$.
We will use the following result which is a corollary of Theorem 3.13 in [26] by Yeo.
Theorem 3.3.
Let $D$ be a semicomplete multipartite digraph. In polynomial time, one can find a cycle factor $F$ with $t$ cycles of $D$, and if $t\geq 2$, an ordering $C_{1},C_{2},\ldots,\\ C_{t}$ of the cycles of $F$ which satisfies the following property: $C_{i}\rightsquigarrow C_{j}$ for all $1\leq i<j\leq t$.
For a cycle $C=x_{1}x_{2}\dots x_{p}x_{1}$ and $i,j\in[p]$, we write $C[x_{i},x_{j}]$ to denote the path $x_{i}x_{i+1}\dots x_{j}\pmod{p}$. By the existence of the ordered cycle factor described in Theorem 3.3, we may obtain the existence of a Hamilton path with special end-vertices in a semicomplete multipartite digraph.
Theorem 1.8. Let $D$ be a semicomplete multipartite digraph and let ${F}$ be a 1-path-cycle factor in $D$. Let $P$ denote the path in ${F}$ and assume the initial and terminal vertices in $P$ belong to different partite sets. Then $D$ contains a Hamilton path whose initial and terminal vertices belong to different partite sets.
Proof.
Let $D$, ${F}$ and $P$ be defined as in the statement of the theorem. Let $P=p_{1}p_{2}p_{3}\ldots p_{\ell}$ and let $D^{\prime}$ be obtained from $D$ by adding the arc $p_{\ell}p_{1}$ if it does not exist in $D$ already (if $p_{\ell}p_{1}\in A(D)$ then $D^{\prime}=D$). Note that $D^{\prime}$ has a cycle factor and let ${C}$ denote a cycle factor with the minimum number of cycles in $D^{\prime}$. If the arc $p_{\ell}p_{1}$ belongs to a cycle in ${C}$ then let $a=p_{\ell}p_{1}$ and otherwise let $a$ be an arbitrary arc on a cycle in ${C}$.
If ${C}$ contains only one cycle then we obtain the desired Hamilton path by removing the arc $a$ from the cycle. So we may assume that ${C}$ contains at least two cycles. We can therefore use Theorem 3.3 and order the cycles, $C_{1},C_{2},\ldots,C_{t}$, of ${C}$ such that $C_{i}\rightsquigarrow C_{j}$ for all $1\leq i<j\leq t$, where $t\geq 2$. Assume that the arc $a$ belongs to $C_{r}$, where $1\leq r\leq t$ and consider the $1$-path-cycle factor, ${F}^{\prime}$, in $D$ obtained from ${C}$ by deleting the arc $a$. Let $P^{\prime}=p_{1}^{\prime}p_{2}^{\prime}p_{3}^{\prime}\ldots p_{m}^{\prime}$ be the path in ${F}^{\prime}$.
Note that $p_{1}^{\prime}$ and $p_{m}^{\prime}$ belong to different partite sets, and assume without loss of generality that $p_{1}^{\prime}\in V_{1}$ and $p_{m}^{\prime}\in V_{2}$, where $V_{1},V_{2},\ldots,V_{c}$ are the partite sets of $D$.
First consider the case when $r<t$. We now transform $P^{\prime}$ and $C_{r+1}$ into a path where the end-points belong to different partite sets as follows. We first consider the case when there is no arc from $C_{r+1}$ to $p_{m}^{\prime}$ in $D$. In this case let $y\in V(C_{r+1})\cap V_{1}$ if such a vertex exists and otherwise let $y\in V(C_{r+1})\setminus V_{2}$ be arbitrary. Note that $p_{m}^{\prime}y\in A(D)$ and the path $P^{\prime}C_{r+1}[y,y^{-}]$ is a path with vertex set $V(C_{r})\cup V(C_{r+1})$ and with end-points in different partite sets.
Secondly we consider the case when there is an arc from $C_{r+1}$ to $p_{m}^{\prime}$ in $D$, say $zp_{m}^{\prime}$. By Theorem 3.3 we note that $p_{m-1}^{\prime}$ and $z^{+}$ both belong to the same partite set $V_{i}$ and the following are both paths on the vertex set $V(C_{r})\cup V(C_{r+1})$.
$P_{1}=p_{1}^{\prime}p_{2}^{\prime}\cdots p_{m-1}^{\prime}zp_{m}C_{r+1}[z^{+},z^{-}]\mbox{ and }P_{2}=p_{1}^{\prime}p_{2}^{\prime}\cdots p_{m}^{\prime}C_{r+1}[z^{+},z]$
And as $z$ and $z^{-}$ belong to different partite sets either $P_{1}$ or $P_{2}$ has initial and terminal vertices from different partite sets. So we have in all cases found a path with vertex set $V(C_{r})\cup V(C_{r+1})$ with initial and terminal vertices in different partite sets.
We can repeat this process in order to get a path with vertex set $V(C_{r})\cup V(C_{r+1})\cup V(C_{r+2})$ with initial and terminal vertices in different partite sets. And continuing this process further we obtain a path with vertex set $V(C_{r})\cup V(C_{r+1})\cup\cdots\cup V(C_{t})$ with initial and terminal vertices in different partite sets.
We can now analogously merge cycles $C_{r-1}$, $C_{r-2}$, $\ldots$, $C_{1}$ with the path ending up with a Hamilton path in $D$ where the initial and terminal vertices are in different partite sets. ∎
Now we characterize the oriented discrepancy of semicomplete multipartite digraphs by the cost of cycle factors in its symmetric (0,1)-digraph.
Theorem 1.7. Let $D$ be an $n$-vertex semicomplete multipartite digraph satisfying the HC-majority inequality and let $c^{cf}_{\max}$ be the maximum cost of a cycle factor in $\hat{D}.$ Let $\sigma^{hc}_{\max}$ be the maximum $\sigma_{\max}(C)$ for a Hamilton oriented cycle $C$ in $D$. Then $\sigma^{hc}_{\max}=c^{cf}_{\max}$ unless $c^{cf}_{\max}=n$ and $D$ is not hamiltonian, in which case $\sigma^{hc}_{\max}=n-1.$ Both $\sigma^{hc}_{\max}$ and a Hamilton oriented cycle $C$ with $\sigma_{\max}(C)=\sigma^{hc}_{\max}$ can be found in polynomial time.
Proof.
Note that a Hamilton cycle $C=x_{1}x_{2}\dots x_{n}x_{1}$ in $\hat{D}$ can be transformed into a Hamilton oriented cycle of $D$ by replacing every arc $x_{i}x_{i+1}\pmod{n}$ of zero-cost by $x_{i+1}x_{i}\pmod{n}$. Also by the construction of $\hat{D}$, $\sigma^{hc}_{\max}$ is equal to the maximum cost of a Hamilton cycle in $\hat{D}.$
Observe that a Hamilton cycle of $\hat{D}$ is a cycle factor of $\hat{D}$. Thus, the maximum cost of a Hamilton cycle of $\hat{D}$ is no more than $c^{cf}_{\max}.$ Let $F$ be a cycle factor of $\hat{D}$ of maximum cost, which can be found in polynomial time, see the last two paragraphs of Theorem 1.6. Let $D_{F}$ be a spanning subdigraph of $\hat{D}$ with $A(D_{F})=A(D)\cup A(F).$ We now consider the following cases.
Case 1: $c^{cf}_{\max}<n$. Thus, $F$ has an arc, $a$, of cost 0. Note that $F-a$ is a 1-path-cycle factor in the semicomplete multipartite digraph $D_{F}-a$. By Theorem 1.8 we note that $D_{F}-a$ contains a Hamilton path, $P=p_{1}p_{2}p_{3}\cdots p_{n}$, with initial and terminal vertices from different partite sets. Now for the Hamilton oriented cycle $C^{\prime}=p_{1}p_{2}p_{3}\cdots p_{n}p_{1}$ in $D$, we know that $\sigma_{\max}(C^{\prime})$ is at least the cost of $P$ and this is at least the cost of $F$ (as the number of zero-cost arcs in $D_{F}-a$ is one smaller than the number of zero-cost arcs in $F$). This implies that $\sigma^{hc}_{\max}=c^{cf}_{\max}$ and completes the proof in this case.
Case 2: $c^{cf}_{\max}=n$. Then $F$ is a cycle factor of $D$. If $D$ is hamiltonian then clearly $\sigma^{hc}_{\max}=n=c^{cf}_{\max}$. So we may assume that $D$ is not hamiltonian. But removing an arc from $F$ and using Theorem 1.8 on the resulting $1$-path-cycle factor in $D$ gives us a Hamilton path in $D$ where the initial and terminal vertices are in different partite sets. We therefore obtain a Hamilton oriented cycle in $D$ with at most 1 backward arc. As $D$ is not hamiltonian, we therefore get $\sigma^{hc}_{\max}=n-1$, which completes the proof in this case.
The proofs of this theorem and Theorem 1.8 can be converted to corresponding polynomial-time algorithms. These algorithms and the polynomial-time algorithms of Theorems 3.2 and 3.3 imply that both $\sigma^{hc}_{\max}$ and a Hamilton oriented cycle $C$ with $\sigma_{\max}(C)=\sigma^{hc}_{\max}$ can be found in polynomial time. ∎
4 Hamilton oriented cycles with maximum number of forward arcs in locally semicomplete digraphs
Let $D$ be a digraph. For disjoint subsets $A$ and $B$ of $V(D)$, if all arcs exist from $A$ to $B$, we say $A$ dominates $B$. We will use the following three results.
Theorem 4.1.
[4] Let $D$ be a connected locally semicomplete digraph that is not strong. Then the strong components of $D$ can be ordered uniquely as $C_{1},C_{2},\dots,C_{\ell}$ such that there is no arc from $V(C_{j})$ to $V(C_{i})$ when $j>i$, $V(C_{i})$ dominates $V(C_{i+1})$ for each $i\in[\ell-1]$ and $C_{t}$ is a semicomplete digraph for each $t\in[\ell]$. And if there is an arc from $V(C_{i})$ to $V(C_{k})$, then $V(C_{i})$ dominates $V(C_{j})$ for all $j=i+1,i+2,\dots,k$ and $V(C_{t})$ dominates $V(C_{k})$ for all $t=i,i+1,\dots,k-1$.
Theorem 4.2.
[4] A connected locally semicomplete digraph has a Hamilton path.
Theorem 4.3.
[4] A strong locally semicomplete digraph has a Hamilton cycle.
Note that an acyclic ordering of strong components in a digraph can be obtained in polynomial time [6]. Thus, the ordering in Theorem 4.1 can be obtained in polynomial time. Using this ordering, it is not hard to construct a Hamilton path in a connected local semicomplete digraph [4]. The proof of Theorem 4.3 in [4] leads to a polynomial-time algorithm for finding a Hamilton cycle in a strong locally semicomplete digraph. Finally, for vertex-disjoint subgraphs $S$ and $T$ of a digraph $D$, it is easy to see that we can find a shortest $(S,T)$-path in polynomial time.
Theorem 4.4.
Let $D$ be a connected locally semicomplete digraph on $n\geq 3$ vertices that is not strong. Let $C_{1},C_{2},\dots,C_{\ell}$ be the acyclic ordering of strong components of $D$. If $U(D)$ is 2-connected, then there is a shortest $(C_{1},C_{\ell})$-path $P$ and a Hamilton oriented cycle $R$ in $D$, such that $P$ is a subpath of $R$ and the set of backward arcs in $R$ is $A(P)$. Such a Hamilton oriented cycle $R$ can be constructed in polynomial time.
Proof.
Let $D$ be a locally semicomplete digraph on $n\geq 3$ vertices that is not strong, but such that $U(D)$ is 2-connected. Let $C_{1},C_{2},\dots,C_{\ell}$ be the acyclic ordering of strong components of $D$ and define the function $cn$ such that $cn(x)=r$ if $x\in V(C_{r})$. Define the $(C_{1},C_{\ell})$-path $P=p_{1}p_{2}\cdots p_{q}$ as follows. Let $p_{1}\in V(C_{1})$ be arbitrary and for each $i$ let $p_{i+1}$ be an arbitrary vertex such that $p_{i}p_{i+1}\in A(D)$ and $cn(p_{i+1})$ is maximum possible. Continue this process until $p_{q}\in C_{\ell}$. By Theorem 4.1 we note that $P$ has length $d(C_{1},C_{\ell})$, i.e., it is a shortest $(C_{1},C_{\ell})$-path in $D$.
We will now show that $D^{\prime}=D-\{p_{2},p_{3},\ldots,p_{q-1}\}$ is connected. For the sake of contradiction assume this is not the case and let $y$ be a vertex which can not be reached from $p_{1}$ in $U[D^{\prime}]$ and such that $cn(y)$ is minimum. This implies that $V(C_{cn(y)-1})=\{p_{i}\}$ for some $i\in\{2,3,\dots,q-1\}$, as otherwise there is a vertex in $V(D^{\prime})\cap V(C_{cn(y)-1})$ which dominates $y$ in $D^{\prime}$, a contradiction to our choice of $y$. As $U[D]$ is $2$-connected $p_{i}$ is not a cut vertex in $U[D]$, which by Theorem 4.1 implies that $V(C_{cn(y)-2})$ dominates $V(C_{cn(y)})$. By our choice of $y$ this implies that $V(C_{cn(y)-2})=\{p_{i-1}\}$. However, this contradicts our construction of $P$ as $p_{i-1}$ has an arc to $y$ and $cn(y)>cn(p_{i})$. Therefore $D^{\prime}$ is connected.
As $D^{\prime}$ is connected, Theorem 4.2 implies that $D^{\prime}$ contains a Hamilton path $Q$. As $Q$ first picks up all vertices in $C_{1}$ and $C_{1}$ is a strong semicomplete digraph and therefore contains a Hamilton cycle, or is a single vertex, we may assume that $Q$ starts in $p_{1}$ (as every vertex in $C_{1}$ have the same out-neighbours in $D-V(C_{1})$). Analogously we may assume that $Q$ ends in $p_{q}$. Now $R=p_{1}Qp_{q}p_{q-1}\cdots p_{1}$ is the desired oriented Hamilton cycle where all arcs on $P$ are backward arcs and all arcs on $Q$ are forward arcs.
Note that using the complexity remarks given after Theorem 4.3, the proof of this theorem can be converted into a polynomial-time algorithm for constructing $R.$
Theorem 4.5.
Let $D$ be a connected locally semicomplete digraph that is not strong. Let $C_{1},C_{2},\dots,C_{\ell}$ be the acyclic ordering of strong components of $D$. For any oriented $(C_{1},C_{\ell})$-path $P$, there is at least $d(C_{1},C_{\ell})$ forward arcs in $P$.
Proof.
Let $D$ and $C_{1},C_{2},\dots,C_{\ell}$ be defined as in the theorem and furthermore define the function $cn$ such that $cn(x)=r$ if $x\in V(C_{r})$. Let $P=p_{1}p_{2}p_{3}\cdots{p_{t}}$ be an oriented $(C_{1},C_{\ell})$-path in $D$. Let $j_{1}=2$. The arc $p_{j_{1}-1}p_{j_{1}}=p_{1}p_{2}$ is a forward arc in $P$ since $P$ is a $(C_{1},C_{l})$-path and vertices in $C_{1}$ do not have in-neighbours in $V(D)\setminus V(C_{1})$. For $k\geq 2$, let $j_{k}$ be the smallest subscript greater than $j_{k-1}$ such that $cn(p_{j_{k-1}})<cn(p_{j_{k}})$ (such $j_{k}$ exists if $j_{k-1}\neq t$ since then $cn(p_{j_{k-1}})<cn(p_{t})$). Observe, from how we choose $j_{k}$, that $cn(p_{j_{k}-1})\leq cn(p_{j_{k-1}})<cn(p_{j_{k}})$. Continue this process and assume that we end up obtaining a sequence $j_{1},j_{2},\dots,j_{q}$ for some positive integer $q$ where $j_{q}=t$. Since $cn(p_{j_{k}-1})\leq cn(p_{j_{k-1}})<cn(p_{j_{k}})$, by Theorem 4.1, we have that for all $k=2,3,4,\dots,q$, $p_{j_{k}-1}p_{j_{k}}\in A(P)$ which in turn gives us $p_{j_{k-1}}p_{j_{k}}\in A(D)$. Thus, $p_{1}p_{j_{1}}p_{j_{2}}\cdots p_{j_{q}}$ is a $(C_{1},C_{\ell})$-path in $D$ with $q$ arcs and therefore $q\geq d(C_{1},C_{\ell})$. In addition, as $p_{j_{k}-1}p_{j_{k}}\in A(P)$ for all $k=1,2,\dots,q$, there are at least $q$ forward arcs in $P$, which combining with $q\geq d(C_{1},C_{\ell})$ proves the result. ∎
Theorem 1.9. Let $D$ be a connected locally semicomplete digraph on $n\geq 3$ vertices. Let $C_{1},C_{2},\dots,C_{\ell}$ be the acyclic ordering of strong components of $D$. Let $\sigma^{hc}_{\max}$ be the maximum $\sigma_{\max}(R)$ for a Hamilton oriented cycle $R$ in $D$. The following now holds.
  • If $D$ is not strong and $U(D)$ is 2-connected, then $\sigma^{hc}_{\max}=n-d(C_{1},C_{\ell})$.
  • If $D$ is not strong and $U(D)$ is not 2-connected, then $D$ contains no oriented Hamilton cycle.
  • If $D$ is strong then $\sigma^{hc}_{\max}=n$.
Furthermore, in polynomial time, we can find a Hamilton oriented cycle of $D$ with the maximum number of forward arcs.
Proof.
We first consider the case when $D$ is not strong and $U(D)$ is 2-connected. Theorem 4.4 implies that $D$ contains an oriented Hamilton cycle $R^{\prime}$, which consists of two internally disjoint paths from a vertex in $C_{1}$ to a vertex in $C_{\ell}$ such that one of them is a shortest $(C_{1},C_{\ell})$-path in $D$. Therefore, $\sigma^{hc}_{\max}\geq\sigma_{\max}(R^{\prime})=n-d(C_{1},C_{\ell})$.
Now let $R=c_{1}c_{2}\cdots c_{n}c_{1}$ be an arbitrary oriented Hamilton cycle in $D$ with the maximum possible number of forward arcs. Without loss of generality, we may assume that $R[c_{a},c_{b}]$ is a $(C_{\ell},C_{1})$-path in $U(D)$, for some $a,b\in[n]$. By Theorem 4.5, we note that the reverse of $R[c_{a},c_{b}]$ contains at least $d(C_{1},C_{\ell})$ forward arcs and therefore $R[c_{a},c_{b}]$ contains at least $d(C_{1},C_{\ell})$ backward arcs. This implies that $R$ contains at least $d(C_{1},C_{\ell})$ backward arcs, which by our choice of $R$ implies that $\sigma^{hc}_{\max}=\sigma_{\max}(R)\leq n-d(C_{1},C_{\ell})$. So, $\sigma^{hc}_{\max}=n-d(C_{1},C_{\ell})$, which completes the first case.
When $D$ is not strong and $U(D)$ is not 2-connected it is clear that $D$ contains no oriented Hamilton cycle, as deleting a vertex from a Hamilton cycle leaves the remaining graph connected. So we may now consider the case when $D$ is strong. However as every strong locally semicomplete digraph contains a Hamilton cycle by Theorem 4.3 we note that this case also holds. Finally, the polynomial-time algorithm claims after Theorem 4.3 and in Theorem 4.4 implies the polynomial-time algorithm claim of this theorem. ∎
5 Complexity of finding Hamilton oriented cycles with maximum number of forward arcs
Encouraged by our results on semicomplete multipartite digraphs and locally semicomplete digraphs, one may guess that if the Hamilton cycle problem is polynomial-time solvable in a class of generalizations of semicomplete digraphs, then the same is true for the problem of finding a Hamilton oriented cycle with maximum number of forward arcs. Unfortunately, this is not the case.
For example, consider locally out-semicomplete digraphs i.e. digraphs in which the out-neighbors of every vertex induce a semicomplete subdigraph [5]. Consider a locally out-semicomplete digraph $H$ obtained from a bipartite undirected graph $B$ with partite sets $X$ and $Y$, $|X|=|Y|\geq 2$, by orienting every edge of $B$ from $X$ to $Y$ and adding an arbitrary arc between every pair of vertices in $Y$. Clearly, $H$ has a Hamilton oriented cycle if and only if $B$ has a Hamilton cycle. However, the problem of deciding whether a bipartite graph has a Hamilton cycle is NP-hard and so is the problem of deciding whether $H$ has a Hamilton oriented cycle.
A digraph $D=(V,A)$ is quasi-transitive if whenever $xy,yz\in A$ for distinct $x,y,z\in V$, we have either $xz\in A$ or $zx\in A$ or both. We can decide whether a quasi-transitive digraph has a Hamilton cycle in polynomial time [21]. However, by orienting all edges of a bipartite graph $B$ from $X$ and $Y$ as above, we obtain a quasi-transitive digraph $Q$, which is even a transitive digraph. Thus, even the problem of deciding whether $Q$ has a Hamilton oriented cycle is NP-hard.
In fact, the above argument shows that it is NP-hard to find a Hamilton oriented cycle in a digraph which is bipartite, acyclic and transitive.
It is natural to ask whether there is a digraph class $\cal D$ such that for $D\in{\cal D}$, in polynomial time, we can decide whether $D$ is hamiltonian and whether $U(D)$ is hamiltonian, but finding a Hamilton oriented cycle with maximum number of forward arcs is NP-hard.
Now consider oriented graphs of Theorem 1.2, i.e., oriented graphs on $n\geq 3$ vertices in which $d(u)\geq n/2$ for every vertex $u$. Is it NP-hard to find a Hamilton oriented cycle with maximum number of forward arcs in this class of oriented graphs? The same question can be asked for oriented graphs of Conjecture 1.3.